OpenStudy (anonymous):

Determine which of the given vectors are not orthogonal: a. u = 3i, v = 8i + 7j b. u = 3j, v = 7i c. u = 3i, v = 7j d. u = 3i – 8j, v = 8i + 3j

OpenStudy (unklerhaukus):

the dot product of orthogonal vectors equals zero

OpenStudy (anonymous):

do they all equal zero?

OpenStudy (unklerhaukus):

only the orthogonal pairs of vectors will have a zero dot product

OpenStudy (unklerhaukus):

what do you get for the dot product of the first pair? u = 3i, v = 8i + 7j u⋅v =

OpenStudy (anonymous):

77

OpenStudy (anonymous):

no wait i got 31

OpenStudy (unklerhaukus):

to compute the dot product of vectors, simply add the products of corresponding components x = x_1i + x_2j y = y_1i + y_2j x⋅y = x_1y_1 + x_2y_2

OpenStudy (anonymous):

was I wrong?

OpenStudy (unklerhaukus):

u = 3i v = 8i + 7j = 3i+0j u⋅v = 3*8 + 0*7 =

OpenStudy (anonymous):

ohhhhhhhhhhhhh i didnt add the 0s...

OpenStudy (anonymous):

the only one not is a!!!

OpenStudy (anonymous):

right?

OpenStudy (unklerhaukus):

that is true but what is a.'s dot product equal to ?

OpenStudy (anonymous):

31

OpenStudy (unklerhaukus):

u⋅v = 3*8 + 0*7 =

OpenStudy (anonymous):

ohhhhh 24

OpenStudy (anonymous):

got it

OpenStudy (unklerhaukus):

yeah 24

OpenStudy (anonymous):

THANK YOU

OpenStudy (unklerhaukus):

now b. u = 3j, v = 7i u⋅v =

OpenStudy (anonymous):

0

OpenStudy (unklerhaukus):

right because 0*3+7*0=0

OpenStudy (unklerhaukus):

c. u = 3i, v = 7j is the almost the same as b.

OpenStudy (unklerhaukus):

d. u = 3i – 8j, v = 8i + 3j u⋅v = .. .. ..